OpenStudy (anonymous):

Sin graph- please help ! y=5sin3x

OpenStudy (anonymous):

period would equal \[\frac{ 2 \pi }{ 3 }\] but how exactly would i draw this \[\frac{ 2 \pi }{ 3 }\] ? :S

OpenStudy (anonymous):

this is a sin wave with amp of 5. and remember the full period of a sin or cos is 2pi. the period just means how long would it take to complete a cycle. you've already figured out thats in this case it's 2pi/3. so lets start with that

OpenStudy (anonymous):